A 50% solution is added to a pure solution to get 40g of 60% solution/ How much pure solution is used.
Answer by stanbon(75887)

A 50% solution is added to a pure solution to get 40g of 60% solution/ How much pure solution is used.
------
Equation:
active + active = active
1*x + 0.50(40-x) = 0.60*40
--------------------
100x + 50*40-50x = 60*40
50x = 10*40
x = 8g (amt. of pure solution needed)
===============================
Cheers,
Stan H.